# Coupled gas and bubble dynamics at the solidification front

### v1

@@ III.4 Where do bubbles nucleate ?
The distance between the solidification front and either the center or edge of the bubble was quantified in relation to the solidification rate $V_{sf}$ (figure 8). Our findings reveal that approximately 73% of the bubbles nucleated during the experiments emerged on the solidification front. This observation suggests a predominantly heterogeneous nucleation mechanism, facilitated by the presence of a substrate (the ice crystals) and the locally elevated gas concentration close to the solidification front due to gas segregation. Nevertheless, it indicates that 27% of the bubbles nucleate within the liquid phase, predominantly within the region extending up to $100\mu m$ from the solidification front, where gas segregation initiates. The solidification velocity does not affect this proportion, as the percentage of bubbles nucleating on the solidification front consistently remains between 60% and 80% across all tested velocities, without any discernible trend.
We estimate that the predominant proportion of the 27% of bubbles that do not nucleate on the solidification front emerges via homogeneous nucleation, attributable to the lack of particles or alternative substrates within the liquid. Conversely, it is plausible that some of these bubbles may have undergone heterogeneous nucleation due to the presence of impurities within the liquid. By considering exclusively those bubbles that nucleated within the bulk liquid, we measured average nucleation distances $d_{sf}$ of $27\pm 25\mu m$ from the bubble edge nearest to the solidification front. No discernible impact of solidification velocity on the measured nucleation distance was identified, consistent with observations made across the entire set of nucleated bubbles. The nucleation distance measurements for bubbles that do not nucleate on the solidification front, across varying solidification velocities, are detailed in Table 1.
@@
Figure 8: Bubble nucleation distance from the solidification front. (A) Measurement of bubble nucleation distances $d_{sf}$ for different solidification velocities $V_{sf}$. These distances are estimated from the initial appearance of the bubble, with the assumption that the displacement of the bubble from its nucleation point to its first observable position is negligible. Measurements were conducted for four different applied solidification velocities, and lateral dispersion was integrated into the data points to improve visual clarity. (B) Proportion of bubbles that nucleate at the solidification front and within the liquid bulk in relation to solidification velocity $V_{sf}$.
It is frequently asserted in solidification studies that all gas bubbles nucleate heterogeneously either on the solidification front [38, 46] or on particles present within the liquid [34, 47]. Moreover, theoretical analyses propose that nucleation is considerably more probable at the solidification front due to its favorable energetic conditions [37, 48]. However, some empirical observations suggest that while the predominant majority of bubbles nucleate heterogeneously on the front, a minor portion may nucleate in proximity to it without the front serving as a substrate [33]. Such nucleation events are typically interpreted as heterogeneous nucleation occurring on minor impurities [39]. Our empirical findings indicate that under the specified experimental conditions, approximately 30% of the bubbles neither nucleate on the solidification front nor on discernible particles or impurities. Although it remains possible that certain bubbles nucleated heterogeneously on the glass slides confining the samples, or on undetectable impurities, we postulate that a significant proportion might have nucleated homogeneously.
@@ III.5 Critical gas concentration for nucleation
The critical bubble nucleation concentration $C_{n}^{*}$ within our system can be derived from experimental observations, using the bubble nucleation distance and the characteristic nucleation time under varying solidification conditions. Assuming that each nucleation event resets the local conditions and initiates a novel phase of gas segregation at the solidification front, Pohl’s equation (Eq. 1) is used to estimate the gas concentration at the specific time and location of nucleation. The interval required to achieve the critical gas concentration facilitating further nucleation corresponds to the lag time $t_{plateau}$ depicted in figure 4 and figure 5. To accurately assess the critical concentration for bubble nucleation, it is crucial to ascertain the nucleation site of the bubbles to incorporate the corresponding distance from the solidification front, denoted as x, into equation 1. Nevertheless, in predicting a homogeneous nucleation threshold, only those bubbles that nucleate at a distance from the solidification front are considered, predicated on the presumption that bubble generation in the liquid bulk primarily results from homogeneous nucleation.
Table 1 provides a summary of the values of the characteristic nucleation time and the homogeneous nucleation distance of bubbles within the liquid across solidification velocities ranging from 5 to $20\mu m/s$. Several estimations of the characteristic bubble nucleation time were conducted. When exclusively considering the plateaus presented in figure 5, the continuous segregation of gas during the nucleation process is disregarded. Nevertheless, as long as the solidification velocity remains constant, the segregation of gas consistently should adhere to equation 1. Consequently, we measured the bubble nucleation times occurring between plateaus to obtain a range of estimations of the critical nucleation concentrations. We observe that the nucleation distance between the bubbles and the solidification interface seems to be independent of the solidification velocity, in contrast to the characteristic nucleation time between successive bubble nucleation bursts $t_{plateau}$.
@@
Table 1: Experimentally measured characteristic data on bubble nucleation dynamics in our system. This table provides the values of the characteristic nucleation time between successive bubble nucleation bursts $t_{plateau}$, the nucleation time, and the nucleation distance of bubbles from the solidification front, with measurements taken at the middle and bottom of the nucleated bubbles. The critical concentrations for bubble nucleation $C_{n}^{*}$ are estimated using equation 1, using the experimentally obtained values presented in this table.
Equation 1 facilitates the reconstruction of gas segregation profiles in accordance with the experimental conditions, and enables the estimation of the critical concentration necessary for bubble nucleation within our system (Figure 9). Consequently, among the range of solidification velocities explored, the principal characteristics distinguishing the critical nucleation concentration are the gas segregation profile and the characteristic lag time tplateau. Our findings indicate that the critical nucleation concentration for bubble formation $C^{*}$ remains consistent across the three solidification velocities examined, demonstrating nearly identical $C^{*}$ values for $V_{sf}=10\mu m/s$, $V_{sf}=15\mu m/s$, and $V_{sf}=20\mu m/s$.
The estimated critical nucleation concentration for $V_{sf}=5\mu m/s$ deviates slightly from those obtained at alternative solidification velocities, with a value approaching $4.5g/L$. Given that the nucleation distance seems to remain largely unaffected by changes in solidification velocity, this variance might originate from the distinctive lag time $t_{plateau}$ observed at this particular velocity. Importantly, $V_{sf}=5\mu m/s$ denotes the scenario where nucleation events were the least frequent, and bubble growth at the interface was notably pronounced, thereby resulting in limited statistical data. Hence, the critical nucleation concentration estimated under these circumstances may be less reliable than at other velocities. Nevertheless, the estimated value retains the same order of magnitude as those derived from other solidification velocities, which we consider satisfactory given the extensive range of critical nucleation concentrations reported in the literature.
Therefore, for the other three velocities examined, it is inferred that the rate of solidification has an insignificant impact on the critical concentration required for bubble nucleation. Analysis of the entire set of experimental data leads to the estimation of an average value $C^{*}=8.4\pm 3.1g/L$, which exceeds the carbon dioxide solubility limit in water at 25C by approximately 5 to 6 times, and approaches 3 times the solubility limit at 0C.
Under our experimental conditions, a portion of the bubble population nucleates heterogeneously at the solidification front. Following nucleation, the majority of these bubbles expand prior to being eventually engulfed by the front. As a result, a significant fraction of the gas that segregates at the solidification front is consumed through both the heterogeneous nucleation of bubbles and their subsequent growth, regardless of the nucleation pathway involved. These findings indicate that the effective gas concentration at the interface preceding bubble nucleation may decrease below the level predicted by equation 1, suggesting that the true critical bubble nucleation concentration $C_{n}^{*}$ under our experimental conditions might be lower than currently estimated.
@@
It is well established that homogeneous nucleation is thermodynamically less favorable compared to heterogeneous nucleation, often necessitating gas supersaturation’s that exceed 100 times the solubility limit in certain systems [29]. The measurement of the critical nucleation concentration in directional solidification presents notable challenges. Various estimates suggest it ranges from 7 [30] to 20 [31], 40 [40] or 100 times [29] the solubility limit, depending on the solidification conditions. Our findings, therefore, fall within the lower spectrum of the estimated critical nucleation concentrations for directional solidification. Two primary hypotheses may account for this relatively low estimation. The first implicates the confinement of our samples between two glass slides, which may also function as preferential nucleation sites. In observing the largest bubbles at the moment of nucleation, the two-dimensional observational approach precludes the determination of whether one of the slides served as a substrate for heterogeneous nucleation. Furthermore, the phenomenon of bubble nucleation predominantly occurring in bursts suggests that the initial bubble in a sequence may perturb the surrounding liquid, thereby inducing either a localized supersaturation or mechanical perturbations conducive to subsequent nearby nucleation. Although this second hypothesis remains plausible, we were unable to confirm it experimentally.
@@
Figure 9: Gas segregation profile at the solidification front during solidification process and estimation of the critical nucleation concentration in our system. The segregation profiles were derived from equation 1, specifically for the instance of carbon dioxide dissolved in water at an initial concentration $C_{0}=1.4g/L$. The critical concentration estimates were obtained using the average nucleation distances and the characteristic nucleation times for the red data points, in addition to the sum of the characteristic nucleation time $t_{n}$ and the nucleation time for the orange data points. (A) $V_{sf}=5\mu m/s$ (B) and $V_{sf}=20\mu m/s$.
@@ IV Conclusions
In addition to the predominance of heterogeneous nucleation at the solidification front, we found that homogeneous nucleation within the liquid phase occurs frequently. The measured nucleation distances for these bubbles do not exhibit a significant dependence on the solidification velocity. These experimental observations enabled us to propose an estimation for the critical nucleation concentration within our system— a critical value that appears to be unaffected by the solidification velocity. These findings enhance our comprehension of the fundamental processes governing bubble nucleation dynamics during solidification, thus contributing to the improved control over porosity formation in solidified materials.

### v2

@@ III.4 Where do bubbles nucleate ?
The distance between the solidification front and either the center or edge of the bubble was quantified in relation to the solidification rate $V_{sf}$ (Fig. 8). Our findings reveal that approximately 73% of the bubbles nucleated during the experiments emerged on the solidification front. This observation suggests a predominantly heterogeneous nucleation mechanism, facilitated by the presence of a substrate (the ice crystals) and the locally elevated gas concentration close to the solidification front due to gas segregation.
It is frequently asserted in solidification studies that all gas bubbles nucleate heterogeneously either on the solidification front [38, 46] or on particles present within the liquid [34, 47]. Moreover, theoretical analyses propose that nucleation is considerably more probable at the solidification front due to its favorable energetic conditions [37, 48]. Our results are therefore in good agreement with previous studies.
We found that 27% of the bubbles nucleate within the liquid phase, predominantly within the region extending up to $100\penalty 10000\ \mu m$ from the solidification front, where gas segregation initiates. The solidification velocity does not affect this proportion, as the percentage of bubbles nucleating on the solidification front consistently remains between 60% and 80% across all tested velocities, without any discernible trend (Fig. 8B). Some empirical observations suggest that while the predominant majority of bubbles nucleate heterogeneously on the front, a minor portion may nucleate in proximity to it without the front serving as a substrate [33]. Such nucleation events are typically interpreted as heterogeneous nucleation occurring on minor impurities [39].
@@
Figure 8: Bubble nucleation distance from the solidification front. (A) Measurement of bubble nucleation distances $d_{sf}$ for different solidification velocities $V_{sf}$. These distances are estimated from the initial appearance of the bubble, with the assumption that the displacement of the bubble from its nucleation point to its first observable position is negligible. Measurements were conducted for four different applied solidification velocities, and lateral dispersion was integrated into the data points to improve visual clarity. (B) Proportion of bubbles that nucleate at the solidification front and ahead of the front in relation to solidification velocity $V_{sf}$.
@@ III.5 Critical gas concentration for nucleation
The critical bubble nucleation concentration $C_{n}^{*}$ within our system can be derived from experimental observations, using the bubble nucleation distance and the characteristic nucleation time under varying solidification conditions. Assuming that each nucleation event resets the local conditions and initiates a novel phase of gas segregation at the solidification front, Pohl’s equation (Eq. 1) is used to estimate the gas concentration at the specific time and distance to the solidification front (Fig. 9). The interval required to achieve the critical gas concentration facilitating further nucleation corresponds to the lag time $t_{plateau}$ depicted in Fig. 4 and Fig. 5. To accurately assess the critical concentration for bubble nucleation, it is crucial to ascertain the nucleation site of the bubbles to incorporate the corresponding distance from the solidification front, denoted as x, into equation 1.
Under our experimental conditions, 73% of the bubble population nucleates heterogeneously at the solidification front ($x=0$). We therefore focus first on this population of bubbles. Under these conditions (Tab. 1), the critical concentration estimated using Eqn. 1 is found to range from $C_{n}^{*}=8.2\pm 0.1\penalty 10000\ g/L$ for $V_{sf}=5\penalty 10000\ \mu m/s$ to $C_{n}^{*}=13.3\pm 0.7\penalty 10000\ g/L$ for $V_{sf}=20\penalty 10000\ \mu m/s$.
Following nucleation, the majority of these bubbles expand prior to being eventually engulfed by the front. As a result, a significant fraction of the gas that segregates at the solidification front is consumed through both the heterogeneous nucleation of bubbles and their subsequent growth, regardless of the nucleation pathway involved. These findings indicate that the effective gas concentration at the interface preceding bubble nucleation may decrease below the level predicted by equation 1, suggesting that the true critical bubble nucleation concentration $C_{n}^{*}$ under our experimental conditions might be lower than currently estimated.
$V_{sf}$ $t_{plateau}$ $t_{plateau}$ nucleation distance x $C_{n}^{*}$ (g/L)
($\mu m/s$) (s) + nucleation time (s) to the front ($\mu m$)
5 $380\pm 130$ $393\pm 132$ 0 $8.2\pm 0.1$
10 $131\pm 43$ $141\pm 53$ 0 $11.1\pm 0.4$
15 $58\pm 23$ $64\pm 26$ 0 $11.2\pm 0.6$
20 $0\pm 11 44$$4\pm 13$ 0 $13.3\pm 0.7$
Table 1: Estimation of the critical concentration for bubble nucleation for bubbles nucleating on the front. This table provides the values of the characteristic nucleation time between successive bubble nucleation bursts $t_{plateau}$, the nucleation time, and the nucleation distance of bubbles from the solidification front (x=0 when nucleation occurs on the front). The critical concentrations for bubble nucleation $C_{n}^{*}$ are estimated using equation 1, using the experimentally obtained values presented in this table.
We focus now on the bubbles nucleating ahead of the solidification front. Among the range of solidification velocities explored, the principal characteristics distinguishing the critical nucleation concentration are the gas segregation profile and the characteristic lag time tplateau. The estimation are provided in Tab. 2. Our findings indicate that the critical nucleation concentration for bubble formation $C_{n}^{*}$ remains consistent across the three solidification velocities examined, demonstrating nearly identical $C_{n}^{*}$ values for $V_{sf}=10\penalty 10000\ \mu m/s$, $V_{sf}=15\penalty 10000\ \mu m/s$, and $V_{sf}=20\penalty 10000\ \mu m/s$. The estimated critical nucleation concentration for $V_{sf}=5\penalty 10000\ \mu m/s$ deviates slightly from those obtained at alternative solidification velocities, with a value approaching $4.5\penalty 10000\ g/L$. Given that the nucleation distance seems to remain largely unaffected by changes in solidification velocity, this variance might originate from the distinctive lag time $t_{plateau}$ observed at this particular velocity. Importantly, $V_{sf}=5\penalty 10000\ \mu m/s$ denotes the scenario where nucleation events were the least frequent, and bubble growth at the interface was notably pronounced, thereby resulting in limited statistical data. Hence, the critical nucleation concentration estimated under these circumstances may be less reliable than at other velocities. Nevertheless, the estimated value retains the same order of magnitude as those derived from other solidification velocities, which we consider satisfactory given the extensive range of critical nucleation concentrations reported in the literature.
Therefore, for the other three velocities examined, it is inferred that the rate of solidification has an insignificant impact on the critical concentration required for bubble nucleation. Analysis of the entire set of experimental data leads to the estimation of an average value $C_{n}^{*}=8.4\pm 3.1\penalty 10000\ g/L$, which exceeds the carbon dioxide solubility limit in water at 25 C by approximately 5 to 6 times, and approaches 3 times the solubility limit at 0 C.
@@
Table 2: Estimation of the critical concentration for bubble nucleation for bubbles nucleating ahead of the front. Experimentally measured characteristic data on bubble nucleation dynamics in our system. This table provides the values of the characteristic nucleation time between successive bubble nucleation bursts $t_{plateau}$, the nucleation time, and the nucleation distance of bubbles from the solidification front, with measurements taken at the middle and bottom of the nucleated bubbles. The critical concentrations for bubble nucleation $C_{n}^{*}$ are estimated using equation 1, using the experimentally obtained values presented in this table.
@@
Regardless of the location of bubble nucleation, on the front or ahead of the front, our results therefore provide a similar estimate of the critical gas concentration for nucleation. The measurement of the critical nucleation concentration in directional solidification presents notable challenges. Various estimates suggest it ranges from 7 [30] to 20 [31], 40 [40] or 100 times [29] the solubility limit, depending on the solidification conditions. Our findings ($\approx$ 3 times the solubility limit), therefore, fall within the lower spectrum of the estimated critical nucleation concentrations for directional solidification. The phenomenon of bubble nucleation predominantly occurring in bursts suggests that the initial bubble in a sequence may perturb the surrounding liquid, thereby inducing either a localized supersaturation or mechanical perturbations conducive to subsequent nearby nucleation. Although this hypothesis remains plausible, we were unable to confirm it experimentally.
@@
Figure 9: Gas segregation profile at the solidification front during solidification process and estimation of the critical nucleation concentration in our system. The segregation profiles were derived from equation 1, specifically for the instance of carbon dioxide dissolved in water at an initial concentration $C_{0}=1.4g/L$. The critical concentration estimates were obtained using the average nucleation distances and the characteristic nucleation times for the red dashed lines, in addition to the sum of the characteristic nucleation time $t_{n}$ and the nucleation time for the orange dashed lines. (A) $V_{sf}=5\mu m/s$ (B) and $V_{sf}=20\mu m/s$.
@@ IV Conclusions
Nucleation was found to be predominantly heterogeneous, occurring on the solidification front. These experimental observations enabled us to propose an estimation for the critical gas concentration for bubble nucleation within our system— a critical value that appears to be unaffected by the solidification velocity. These findings enhance our comprehension of the fundamental processes governing bubble nucleation dynamics during solidification, thus contributing to the improved control over porosity formation in solidified materials.